OpenStudy (anonymous):

Find the volume of each cylinder. Give your answers both in terms of  and rounded to the nearest tenth.

OpenStudy (anonymous):

A cylinder with diameter 20 inches and height 2in

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The diameter is 20. d = 20

jimthompson5910 (jim_thompson5910):

the radius is half of the diameter r = d/2 r = 20/2 r = 10

jimthompson5910 (jim_thompson5910):

now use the formula \[\Large V = \pi*r^2*h\]

jimthompson5910 (jim_thompson5910):

r = radius h = height so, r = 10 h = 2

OpenStudy (anonymous):

Omg thank you so much

OpenStudy (anonymous):

You deserve a medal

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (anonymous):

so is the answer 100 x 2?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's 200 and don't forget the pi

jimthompson5910 (jim_thompson5910):

your exact volume in terms of pi is \[\Large V = 200\pi\]

jimthompson5910 (jim_thompson5910):

then use \[\Large \pi \approx 3.14\] to get the approximate volume

jimthompson5910 (jim_thompson5910):

use more digits of pi to get a more accurate approximation

OpenStudy (anonymous):

Okay thank you . You are so awesome

jimthompson5910 (jim_thompson5910):

np